OpenStudy (anonymous):

z/2=y/x.....In the equation above, the numbers x, y, and z are either positive or negative. Which of the following must be equal to 4? a.2x b.2z c.xz/2y d.2y/xz e.2xz/y

OpenStudy (luigi0210):

\[\frac{ z }{ 2 }=\frac{ y }{ x }\] \[zx=2y\] \[4=\frac{ 2y }{ zx }\] total guess.. should probably ask someone else :/

OpenStudy (anonymous):

thanks anyway

OpenStudy (valpey):

Cross multiply and you see that 2y = zx Therefore 2 = zx/y Multiply both sides by 2.

OpenStudy (valpey):

(important that we know neither y or x is zero)